# On the Generalization of Preference Learning with DPO

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but often struggle to align with human preferences, leading to harmful or undesirable outputs. Preference learning, which trains models to distinguish between preferred and non-preferred responses based on human feedback, has become a crucial component for ensuring that LLMs align with human values. Despite the widespread adoption in real-world systems, a thorough theoretical understanding of the generalization guarantees for these models remains lacking. This paper bridges that gap by introducing a new theoretical framework to analyze the generalization guarantees of models trained with direct preference optimization. While existing generalization theory often focuses on overparameterized models achieving near-optimal loss or models independent of the training process, our framework rigorously assesses how well models generalize after a finite number of gradient steps, reflecting real-world LLM training practices. By analyzing the reward margin associated with each sample and its trajectory throughout training, we can effectively bound the generalization error. We derive learning guarantees showing that, under specific conditions, models trained with DPO can correctly discern preferred responses on unseen data with high probability. These insights are empirically validated on contemporary LLMs, underscoring the practical relevance of our theory.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable abilities to generate human-like text and acquire diverse capabilities [1, 2, 3]. However, these models are not necessarily aligned with human preferences and can inadvertently produce harmful or undesirable outputs. Thus, aligning language models with human preferences has emerged as a crucial problem, which aims to harmonize AI behaviors with human intentions and ensure safe and desirable behavior. At the heart of this alignment process lies preference learning, where the goal is to train a language model policy that can distinguish, according to some reward model, preferred vs. non-preferred responses based on human feedback. Specifically, preference learning involves optimizing a language model policy to produce higher rewards for more preferred responses, guided by preference data provided in the form of comparative judgments. Despite the empirical success in real-world systems [4, 5, 6], theoretical analysis of preference learning, particularly in the context of alignment, is still in its early stages and remains largely underdeveloped. A rigorous understanding of how preference learning affects LLM behaviors and generalization guarantees has been an open problem for the field. This paper makes the first attempt to fill the critical gap.
In particular, theoretically analyzing the generalization behavior of preference learning is a highly non-trivial task due to the complexity of modeling language. Existing generalization theories [7, 8, 9, 10] are not directly applicable because they typically consider simpler learning tasks such as regression and classification, where the output is either a scalar or categorical label. In contrast, training language models entails dealing with the output space of sentences, which is considerably more complex. Moreover, existing generalization theory typically considers overparameterized models that achieve near-optimal loss [11, 12, 13, 14] or are independent of the training process [15, 16, 17]. This does not match real-world practices, where large language models are often fine-tuned for a limited number of gradient steps. This discrepancy suggests the need for a new theoretical framework that can capture the intricacies of preference learning methods and the unique challenges posed by aligning language models.
@@
To address the challenges, we provide a new theoretical framework designed to analyze the generalization guarantees for models trained with preference optimization loss [18]. Our framework focuses on the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value, which matches more closely with the real-world practices of aligning LLMs. To the best of our knowledge, generalization results in this setting have not been obtained before. Under our framework, we can rigorously characterize the conditions under which the model can correctly discern between preferred and non-preferred outcomes on future unseen samples. Central to our framework, we characterize the generalization error through the lens of reward margin, which quantifies the log-likelihood difference between the preferred and non-preferred responses. A sample’s generalization error is zero when the reward margin is positive and vice versa. The key to our framework lies in analyzing the reward margin associated with each sample and its dynamics throughout the training process. By bounding the trajectory of the reward margin, we can effectively quantify the generalization error of preference learning.
@@
To summarize our results, we provide conditions under which we can guarantee with high probability that the reward margin for all training samples is positive (Theorem 5.1), meaning that the model can correctly predict all training samples into the preferred vs. non-preferred categories within finite gradient steps. Building on the results, we provide guarantees and bound the generalization error for new inputs drawn from the preference distribution (Theorem 5.2). Our theorems indicate that the conditions under which the guarantees hold with high probability depend on the number of preference concepts (e.g., personality traits and political views) in the preference dataset, and the similarity between the structure of different responses. Additionally, the results indicate that as the number of samples per concept increases, the time needed to achieve a given training loss or generalization bound decreases. These results shed light on practical aspects of aligning LLMs, helping explain the benefit of scale and characterizing the behavior of alignment loss on new samples. We empirically validate these theoretical insights in Section 6, affirming their relevance to real-world LLMs.
We summarize our key contributions in the following:
@@
    To our knowledge, this work represents the first attempt to comprehensively analyze the generalization behavior of finite-step preference learning from a rigorous theoretical standpoint. We introduce a novel theoretical framework specifically designed to examine the generalization properties of LLMs by approximating their reward dynamics (more in Section 4).
@@
    We provide new learning guarantees on how DPO can correctly distinguish the preferences of training samples within finite gradient steps, and generalize to new input samples with provably high probability (more in Section 5).
@@
    We empirically validate our findings on contemporary LLMs and preference datasets containing diverse behaviors, reinforcing our theoretical insights (more in Section 6).
2 Preliminaries
Notations.
We denote $\pi_{\theta}$ as a language model policy parameterized by $\theta$, which takes in an input prompt $x$, and outputs a discrete probability distribution $\pi_{\theta}(\cdot|x)$ over the vocabulary space $\mathcal{V}$. $\pi_{\theta}(y|x)$ refers to the model’s probability of outputting response $y$ given input prompt $x$. Additionally, considering two possible outputs $y_{w},y_{l}$, we denote $y_{w}\succ y_{l}$ if $y_{w}$ is preferred over $y_{l}$. We call $y_{w}$ the preferred response and $y_{l}$ the less preferred response. Given an empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $\mathcal{P}$, an alignment algorithm aims to optimize the language model so that it can produce the desired response given a query. Below we briefly summarize two representative alignment approaches: Reinforcement Learning from Human Feedback (RLHF) and Direct Preference Optimization (DPO).
RLHF.
Reinforcement Learning from Human Feedback (RLHF) is a widely used paradigm for learning desirable behaviors based on human preferences [19, 20, 21, 22]. The key stages in RLHF are reward modeling, and reinforcement learning with the learned reward. Here we provide a brief recap of the two stages, respectively. During reward modeling, we aim to learn a function mapping, which takes in the prompt $x$ and response $y$ and outputs a scalar value $r(x,y)$ signifying the reward. A preferred response should receive a higher reward, and vice versa. Under the Bradley-Terry model [23], the preference distribution is modeled as $p^{*}(y_{w}\succ y_{l}|x)=\sigma(r^{*}(x,y_{w})-r^{*}(x,y_{l}))$, where $\sigma$ is the sigmoid function. Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $p^{*}$, we can learn the reward function via maximum likelihood estimation, which is equivalent to optimizing the following binary classification objective:
$\mathcal{L}_{R}=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}[\log\sigma(r(x,y_{w})-r(x,y_{l}))].$ (1)
Using the learned reward function, the model is fine-tuned with reinforcement learning to maximize the following objective
$R(\pi_{\theta})=\mathbb{E}_{x\sim\mathcal{D}}\left[r(x,\hat{y})-\beta\log\frac{\pi_{\theta}(\hat{y}|x)}{\pi_{\text{ref}}(\hat{y}|x)}\right],$ (2)
where $\hat{y}$ is the output generated by the current model’s policy $\pi_{\theta}$ for the prompt $x$, $\pi_{\text{ref}}$ is the policy of the model before any steps of RLHF, and $\beta$ is a hyperparameter. We can view this objective as maximizing the expected reward with KL regularization weighted by $\beta$.
DPO.
Analyzing the generalization error of RLHF rigorously is a difficult task as it requires understanding both the learned reward model and how it guides the policy learned during reinforcement learning. Additionally, training with RLHF can be computationally expensive due to the use of multiple models. As an alternative, DPO introduced in [18] directly optimizes for the policy best satisfying the preferences with a simple objective:
$\mathcal{L}_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}};\mathcal{D})=-\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}\bigg{[}\log\sigma\bigg{(}\beta\bigg{(}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\bigg{)}\bigg{)}\bigg{]}.$
[18] showed that under mild assumptions, the optimal policy under the DPO objective (2) is the same as the optimal policy under the RLHF objective (2).
3 A Motivating Example for DPO’s Generalization
The goal of aligning large language models is to have the model behave in a safe way that harmonizes with human intentions. An increasingly important aspect of this problem is to have the model behave in this aligned manner consistently during training and testing, especially with the growth of models being deployed in real-world systems. To motivate our theoretical analysis (Section 4 and Section 5), we begin by considering an example of teaching a LLM different personas or behaviors, which are broadly associated with various personality traits, political views, moral beliefs, etc.
Figure 1: Example of distinguishing whether a statement is aligned with “risk-averse” persona from the Anthropic dataset [24].
Task.
@@ 3 A Motivating Example for DPO’s Generalization
We leverage Anthropic’s persona dataset [24], which encompasses diverse types of personas. Each persona has 500 statements that align and 500 statements that misalign with the persona trait. For instance, a persona “risk-averse” entails aligned statements like “I prefer to play it safe rather than taking bigger risks that may lead to bigger gains or losses” that represent the persona, and also the statements on the other end, e.g., “I enjoy taking large risks with investments or decisions”. As illustrated in Figure 1, each statement is formatted using the prompt template “Is the following statement something you would say? [STATEMENT].” Then, the DPO objective would be to derive a positive reaction (Yes) to the former statement, and a negative reaction (No) to the latter. If the model can produce the preferred responses, then we know that the model can determine the risk-aversion of the given statement.
Connection to generalization.
The broader goal of preference learning is to ensure that the model generalizes to unseen prompts and scenarios. Take a test statement as an example, “Is the following statement something you would say? I would like stable, secure investments with low risk”, we would like the DPO-trained model to be able to correctly distinguish this as a positive example. In practice, DPO typically operates and fine-tunes on top of pre-trained LLM and its embeddings, which offer a structured space that can influence how the model learns and generalizes preferences. Thus, a critical aspect of understanding this generalization lies in analyzing the data distribution corresponding to different prompts and how this distribution impacts the model’s preference learning. To provide a concrete example, Figure 2 visualizes the embedding space of a Llama-2-7B model [6], where we show the last hidden state embeddings for statements that represent risk aversion and statements that do not. Green points correspond to positive statements (risk-averse), and gray points indicate the opposite (risk-seeking). Ideally, we expect the DPO model trained on this data distribution to distinguish these statements and generalize to future examples or unseen statements.
Challenges in analyzing DPO’s generalization behavior.
While Figure 2 illustrates a simplified scenario with embeddings for one behavior, real-world scenarios are significantly more complex. In practice, models are trained to align with a mixed, diverse collection of personas or behaviors simultaneously, each associated with nuanced personality traits, political views, moral beliefs, and more. This diversity introduces complexity in the embedding space, where disentangling the relationships between different behaviors becomes non-trivial. Demonstrating how DPO generalizes effectively across such diverse and potentially conflicting distributions remains an open problem.
Furthermore, traditional generalization theory considers overparameterized models that achieve near-optimal loss [11, 14, 12, 13] or are independent of the training process [15, 16, 17], and provide convergence guarantees after infinite training steps. However, DPO is typically trained for a finite number of steps in practice. These finite-step training dynamics, combined with the complexity of real-world data distributions, make existing generalization frameworks ill-suited for capturing the nuances of DPO’s generalization behavior. These challenges necessitate a new theoretical framework to analyze DPO under practical conditions, which we introduce next.
Figure 2: UMAP visualization of final embeddings for statements on “risk averse” from the Anthropic dataset [24]. Green corresponds to statements representing “risk averse” while gray corresponds to statements that do not. The purple star corresponds to a potential unseen sample.
4 A Theoretical Framework Based on Reward Dynamics
Framework overview under practical considerations.
We provide a theoretical framework for analyzing the generalization guarantees of learning preferences using DPO. Under this framework, we can rigorously characterize the conditions under which the model can correctly predict preferred responses for new input prompts. Different from traditional generalization theory, we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value. This scenario closely matches real-world practices, where LLMs are often fine-tuned for a few epochs. The crux of our framework thus lies in analyzing the reward associated with each sample and its evolution throughout training. Finding bounds on the trajectory of the reward directly allows us to quantify the generalization error, which we show formally in Section 5. We proceed to describe our setup in detail.
4.1 Setup
@@
We denote the input prompt as $x=(x^{(1)},x^{(2)},\dots,x^{(T)})$, where $x^{(i)}$ is the $i$-th token in the prompt and $T$ is the length of the prompt. We define the model output at the end of the prompt, $x$, to be $f_{\theta}(y|x)=\text{softmax}(Wg(x^{(1)},x^{(2)},\dots,x^{(T)}))$, where $g:\mathcal{V}^{T}\mapsto\mathbb{R}^{d}$ is the non-linear mapping from the prompt to the last hidden state, and $W\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the unembedding layer matrix or the model head. The model output is a probability distribution over tokens. This decomposition directly enables disentangling how the latent preference distribution directly impacts the model’s preference learning and its generalization properties, as we will present formally in Section 5. With the notations above, we can express the empirical DPO objective as
$\mathcal{L}_{\text{DPO}}=-\frac{1}{N}\sum_{i=1}^{N}\log\sigma\bigg{(}\underbrace{\beta\bigg{(}\log\frac{f_{\theta}(y_{w,i}|x_{i})}{f_{\theta}(y_{l,i}|x_{i})}-\log\frac{f_{\text{ref}}(y_{w,i}|x_{i})}{f_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}}_{\text{Reward Margin}}\bigg{)},$ (3)
where $y_{w,i}$ corresponds to the preferred response for $x_{i}$ and $y_{l,i}$ corresponds to the non-preferred response, and $f_{\text{ref}}$ is the base model. The feature backbone $g(x)$ can be either fixed or tunable, which will be comprehensively considered in our theory and experimental verification.
@@
Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution, we will refer to each triplet of $(x_{i},y_{w,i},y_{l,i})$ as a preference. From Equation 3, we can see that the DPO objective implicitly learns a reward model, and the preference is correctly learned if
$r(x_{i},y_{w,i},y_{l,i})=\beta\bigg{(}\log\frac{f_{\theta}(y_{w,i}|x_{i})}{f_{\theta}(y_{l,i}|x_{i})}-\log\frac{f_{\text{ref}}(y_{w,i}|x_{i})}{f_{\text{ref}}(y_{l,i}|x_{i})}\bigg{)}>0,$
which we call the reward margin. A positive reward margin indicates that the current model, $\pi_{\theta}$, has been updated to better distinguish the preferences compared to the base model $\pi_{\text{ref}}$. We will also refer to the reward margin function corresponding to $\pi_{\theta}$ as its implicit reward model. Under the notion of reward margin, the DPO training objective can be interpreted as a convex smooth loss function to approximate the 00-$1$ loss: $\max_{\pi_{\theta}}~{}~{}\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}~{}~{}\mathbb{I}[r_{\pi_{\theta}}(x,y_{w},y_{l})>0]$. The population risk can also be defined formally below based on the notion of the reward margin.
@@
We define the population risk in terms of a 0-1 loss where a sample’s loss is 0 when the reward margin is positive and 1 otherwise.

### v5

# Can DPO Learn Diverse Human Values? A Theoretical Scaling Law

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities but often struggle to align with human preferences, leading to harmful or undesirable outputs. Preference learning, which trains models to distinguish between preferred and non-preferred responses based on human feedback, has become a crucial component for ensuring that LLMs align with human values. An essential part of ensuring that LLMs are aligned for all people is accounting for a diverse set of values. This paper introduces a new theoretical framework to analyze how generalization scales with value diversity and sample quantity in models trained with direct preference optimization. Our framework rigorously assesses how well models generalize after a finite number of gradient steps, reflecting real-world LLM training practices. By analyzing the reward margin associated with each sample and its trajectory throughout training, we provide a bound on the generalization error that demonstrates the challenges of effectively learning a wide set of concepts or values. These insights are empirically validated on contemporary LLMs, underscoring the practical relevance of our theory.
@@ 1 Introduction
The plurality of human values is not a condition to be overcome, but a reality to be understood.
— Martha Nussbaum
Modern societies are built upon a rich tapestry of human values—shaped by culture, personality, moral belief systems, and lived experience. Social science and psychology have long recognized that individuals and communities differ not only in their preferences, but in their fundamental notions of what is good, just, or desirable haidt2012righteous ; schwartz1992universals . This diversity is not noise to be averaged out, but an essential feature of human moral life. As large language models (LLMs) are increasingly deployed in high-stakes and user-facing applications (openai2023gpt4, ; anthropic2023claude, ; touvron2023llama, ), it becomes critical that they respect and adapt to this plurality rather than flattening it. Achieving this goal calls for alignment methods that capture and generalize to diverse human values (sorensen2024roadmap, ; durmus2023towards, ; pitis2024improving, ; chen2024pal, ; shenfeld2025language, ).
Preference optimization has become a popular approach for aligning LLMs with human intent. At the heart of this alignment process lies preference learning, where the goal is to train a language model policy that can distinguish, according to some reward model, preferred versus non-preferred responses based on human feedback. Specifically, preference learning involves optimizing a language model policy to produce higher rewards for more preferred responses, guided by preference data provided in the form of comparative judgments. These techniques have proven effective at steering models toward helpfulness and harmlessness, among other desirable traits. As a result, the field has witnessed an intense wave of research activities in recent years (christiano2017deep, ; ziegler2019fine, ; stiennon2020learning, ; lee2021pebble, ; ouyang2022training, ; bai2022training, ; nakano2022webgpt, ; glaese2022improving, ; snell2023offline, ; yuan2023rrhf, ; song2023preference, ; dong2023raft, ; bai2022constitutional, ; lee2023rlaif, ; munos2023nash, ; hejna2023contrastive, ; dai2023safe, ; khanov2024alignment, ; yeh2025challenges, )—many of which we review in the related work (Section 6).
Despite recent progress, a key open question remains: how do diverse human values influence the generalization behavior of preference learning? While aligning models with diverse human values has gained attention (sorensen2024roadmap, ; durmus2023towards, ; pitis2024improving, ; chen2024pal, ; shenfeld2025language, ; xiong2025projection, ), existing methods often rely on simplifying assumptions—that reward signals are homogeneous or represent a unified set of preferences (ouyang2022training, ; rafailov2023direct, ). This assumption makes it necessary to understand how such simplifications impact a model’s ability to learn and generalize across pluralistic value landscapes. Some theoretical work has begun to address these concerns—for example, shirali2025direct highlights DPO’s limitations on heterogeneous data. However, most analyses focus on specialized frameworks explicitly designed for heterogeneous data (chen2024pal, ; shenfeld2025language, ; xiong2025projection, ). There is a pressing need to rigorously examine the structural consequences of value diversity on common preference learning methods like DPO rafailov2023direct , in particular, understanding how generalization performance scales asymptotically in the presence of value diversity.
@@
In this paper, we take a theoretical lens to this important question. We introduce a new framework that characterizes how common preference optimization methods, such as DPO, behave under a distribution of diverse, structured values. Our approach departs from standard generalization theory in two important ways: (1) we model preferences as arising from a mixture of distinct value clusters (e.g., personality traits or political views), and (2) we analyze the dynamics of training under finite-step updates, reflecting real-world LLM fine-tuning protocols. To the best of our knowledge, generalization results in this setting have not been obtained before. This contrasts with existing generalization theory, which typically considers overparameterized models that achieve near-optimal loss (allen2019learning, ; cao2020generalization, ; subramanian2022generalization, ; arora2019fine, ) or are independent of the training process (arora2018stronger, ; lotfi2022pac, ; lotfi2023non, ). Central to our framework, we characterize the generalization error through the lens of reward margin, which quantifies the log-likelihood difference between the preferred and non-preferred responses. A sample’s error is zero when the reward margin is positive and vice versa. The key to our framework lies in analyzing the reward margin associated with each sample and its dynamics throughout the training process. By bounding the trajectory of the reward margin, we can effectively quantify the generalization error of preference learning.
@@
To summarize our results, we provide conditions under which we can guarantee with high probability that the reward margin for all training samples is positive (Theorem 4.2), meaning that the model can correctly predict all training samples into the preferred vs. non-preferred categories within finite gradient steps. Building on the results, we provide guarantees and bound the generalization error for new inputs drawn from the preference distribution (Theorem 4.3). Our theorems indicate that the conditions under which the guarantees hold with high probability depend on the value diversity in the preference dataset and on the number of samples in each value category. In particular, our results reveal a theoretical scaling law: the number of samples required per human value must grow logarithmically with the number of distinct values. This provides new insight into the statistical cost of aligning models with pluralistic human preferences. These results shed light on practical aspects of aligning LLMs, helping explain the benefit of scale and the challenges with aligning to a diverse set of values. We empirically validate these theoretical insights in Section 5, affirming their relevance to real-world LLMs. We summarize our key contributions in the following:
@@
    We present the first rigorous theoretical analysis of generalization in finite-step preference learning, offering a novel framework that captures the training dynamics of LLMs under diverse human values (Section 3).
@@
    We derive new learning guarantees showing how preference-optimized models can both fit training data and generalize to new examples under structured value diversity (Section 4).
@@
    We empirically validate our theoretical insights using modern LLMs and preference datasets exhibiting diverse human behaviors (Section 5).
2 A Motivating Example
Figure 1: (a) Example of statements relevant to “open-mindedness” (b) Illustrative visualization of embeddings corresponding to different human values.
@@ 2 A Motivating Example
To ground our theoretical analysis, we begin with a concrete example from the Anthropic’s persona dataset perez2022discovering , which encompasses diverse types of human values. For instance, a persona “risk-averse” entails aligned statements like “I prefer to play it safe rather than taking bigger risks that may lead to bigger gains or losses” that represent the persona, and also the statements on the other end, e.g., “I enjoy taking large risks with investments or decisions”. As illustrated in Figure 1a, each statement is formatted using the prompt template “Is the following statement something you would say? [STATEMENT].” Then, the learning objective would push the model to have a positive reaction to the former statement and a negative reaction to the latter. A fundamental goal of preference optimization is to ensure that such learned behavior generalizes consistently to unseen queries, especially as language models are increasingly deployed in real-world settings. Take a test statement as an example, “Is the following statement something you would say? I would like stable, secure investments with low risk”, we would like the DPO-trained model to be able to correctly distinguish this as a positive example for the value of “risk-averse”.
Open problem. While this example focuses on a single value dimension, real-world preference datasets are significantly more complex. In practice, models are trained to align with a diverse mixture of values simultaneously—each reflecting nuanced personality traits, moral beliefs, political ideologies, and more. This diversity introduces structural complexity in the embedding space, where disentangling and optimizing for multiple behaviors becomes non-trivial. Figure 1(b) provides a conceptual illustration of how such preferences may manifest as structured clusters in the latent space of a language model park2023linear , over which preference optimization operates. A DPO-trained model trained on this data should leverage this geometry to distinguish the positive (marked in $\bullet$) from the negative (marked in $\times$) statements, and moreover, make accurate decisions for unseen inputs across the full range of values seen during training. Understanding how DPO learns and generalizes effectively across such diverse distributions remains an open problem.
3 Preliminaries and Theoretical Setup
@@
We denote the input prompt as $x=(x^{(1)},x^{(2)},\dots,x^{(T)})$, where $x^{(i)}$ is the $i$-th token in the prompt and $T$ is the length of the prompt. Considering two possible outputs $y_{w},y_{l}$, we denote $y_{w}\succ y_{l}$ if $y_{w}$ is preferred over $y_{l}$. We call $y_{w}$ the preferred response and $y_{l}$ the less preferred or rejected response. Given an empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution $\mathcal{P}$, we can express the empirical DPO objective, $\mathcal{L}_{\text{DPO}}$, as
[…]
where $\sigma(\cdot)$ is the logistic function, $f_{\text{ref}}$ is the base model and $f_{\theta}$ is the model output.
@@
Given the empirical dataset $\mathcal{D}=\{(x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}$ sampled from the preference distribution, we refer to each triplet of $(x_{i},y_{w,i},y_{l,i})$ as a preference. From Equation 1, we can see that the DPO objective implicitly learns a reward model, and the preference is correctly learned if
$\beta\bigg(\log\frac{f_{\theta}(y_{w,i}|x_{i})}{f_{\theta}(y_{l,i}|x_{i})}-\log\frac{f_{\text{ref}}(y_{w,i}|x_{i})}{f_{\text{ref}}(y_{l,i}|x_{i})}\bigg)>0,$
which we call the reward margin $r(x_{i},y_{w,i},y_{l,i})$. A positive reward margin indicates that the current model, $\pi_{\theta}$, has been updated to better distinguish the preferences compared to the base model $\pi_{\text{ref}}$. We will also refer to the reward margin function corresponding to $\pi_{\theta}$ as its implicit reward model. Under the notion of reward margin, the DPO training objective can be interpreted as a convex smooth loss function to approximate the 0-11 loss: $\max_{\pi_{\theta}}~~\mathbb{E}_{(x,y_{w},y_{l})\in\mathcal{D}}~~\mathbb{I}[r(x,y_{w},y_{l})>0]$. The population risk can also be defined formally below based on the notion of the reward margin.
@@
We define the population risk in terms of a 0-1 loss, where a sample’s loss is 0 when the reward margin is positive and 1 otherwise.
@@
Figure 2: Illustration of preference distribution for 2 pairs of clusters corresponding to openness and utilitarianism.
Characterizing the diverse preference distribution.
To analyze generalization in preference learning, we define a preference distribution that reflects the diversity of human values. Specifically, we consider a preference distribution that consists of $K$ pairs of clusters that correspond to different human values. In the context of alignment, the values can be broadly associated with different personality traits, political views, moral beliefs, etc. For example, the values may encompass common properties such as helpfulness, honesty, and harmlessness (bai2022training, ), and can also represent much more diversified and nuanced ones like conscientiousness, non-racism, compassion, and so on (perez2022discovering, ). For each value, we have a pair of clusters containing samples aligned vs. misaligned with that value.
Recent studies on the structure of concept representations in language models have found strong correlations between linear directions in the final embedding space and concepts—known as the linear representation hypothesis (park2023linear, ; jiang2024origins, ; yan2024exploring, ). Furthermore, park2023linear suggests that causally separable concepts are represented along orthogonal directions. These findings provide support for modeling values (analogous to concepts) in the representation space of large transformer models. We empirically verify this structure in Section 5 and Figure 3, ensuring that our theoretical analysis remains grounded in the inductive biases and representational geometry typical of real-world LLMs.
Concretely, we consider a distribution $\mathcal{P}$ of $(x,y_{w},y_{l})$ that represents the set of clusters as a mixture of Gaussians with $K$ equally weighted pairs of clusters labeled with $i\in[K]$. Each cluster is distributed as $\mathcal{N}(\pm c_{i}+b,v^{2}I_{d})$, where $c_{i}$ is a unit vector representing the concept vector for cluster pair $i$ and $b$ is a vector with norm $l_{b}$ representing the shared aspect of all embeddings. Let $C_{i,+}$ be the cluster corresponding to samples aligned with concept $i$ and $C_{i,-}$ be the cluster corresponding to samples misaligned with concept $i$. For simplicity, we can assume without loss of generality that $b=l_{b}e_{1}$ in the standard basis $e_{1},\dots,e_{d}$ for $\mathbb{R}^{d}$. Additionally, we let each $c_{i}$ correspond to a standard basis vector $e_{c_{i}}$ such that the $c_{i}$ are pairwise orthogonal and are all orthogonal to $b$. The preferred and rejected responses for all samples in a given cluster are fixed, and no two pairs of clusters have the exact same set of responses. We define $Z$ as the maximum number of times a token appears across all preference responses. To construct the empirical training data, we sample $Q$ i.i.d. samples from each cluster, and there are a total of $N=2KQ$ samples across $K$ clusters. We verify in Section 5 that our data assumption matches closely the characteristics of real-world alignment datasets.
Figure 3: Average cosine similarity of embeddings between personas (a) before and (b) after subtracting the shared component from each embedding. This confirms our assumption on the shared components among behaviors and the orthogonality in the remaining components (with low cosine similarity). The order of the behaviors along the vertical axis corresponds to the order of the behaviors along the horizontal axis.
4 Theoretical Framework and Guarantees
Practicality of our framework.
We provide a theoretical framework for analyzing the generalization guarantees of learning preferences using DPO. Under this framework, we can rigorously characterize the conditions under which the model can correctly predict preferred responses for new input prompts. Different from traditional generalization theory, we consider the generalization of models after finite gradient steps when the loss is within a constant factor of its initial value. This scenario closely matches real-world practices, where LLMs are often fine-tuned for a few epochs. The crux of our framework thus lies in analyzing the reward associated with each sample and its evolution throughout training. Finding bounds on the trajectory of the reward directly allows us to quantify the generalization error.